Two people walking past each other. They continue in opposite directions. The 1st person is walking at 6mph and the other one is speed walking at 10mph. How much must they walk to be at least 2 miles apart?
_______minutes (in decimal form)
thanx.

Found 3 solutions by Nate, checkley71, stanbon:
Answer by Nate(3500)

10x + 6x = 2
16x = 2
x = 0.125
.125 * 60 = 7.5 minutes

Answer by checkley71(8403)

6X+10X=2 OR 16X=2 OR X=2/16 OR .125 HOUR OR 3.75 MIN
PROOF 6*.125+10*.125=2 OR .75+1.25=2 OR 2=2

Answer by stanbon(48569)

Two people walking past each other. They continue in opposite directions. The 1st person is walking at 6mph and the other one is speed walking at 10mph. How much must they walk to be at least 2 miles apart?
_______minutes (in decimal form)
Let the time each one walks be "x" hrs.
Then the 1st person will walk 6x miles
The 2nd person will walk 10x miles
INEQUAILITY:
distance 1st + distance 2nd >=2 miles
6x+10x>=2
x>=1/8 hr = 0.125 hr. = (0.125)(60min)=7.5 minutes
Cheers,
Stan H.